OpenStudy (anonymous):

slove for x 13x+7=5x-20

OpenStudy (anonymous):

x = - 27/8

OpenStudy (imtiaz7):

13x-5x=-20-7 8x=-27 x=-27/8

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Anytime.

OpenStudy (imtiaz7):

welcome :)

OpenStudy (anonymous):

Could you guys please help with my question?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

I posed it on my profile. If any of you guys could help, that would be really awesome.